Let B = (a1, a2, a3) be an ordered basis of  R3 with a1 = (1, 0, -1), a2 = (1, 1, 1),  a3 = (1, 0, 0).

Write the vector v = (a, b, c) as  a linear combination of the basis vectors  from B.


1
Expert's answer
2020-11-09T20:10:25-0500

Let "v = (a, b, c)=pa_1+qa_2+ra_3" for some "p,q,r\in\mathbb R" and find the coefficients "p,q,r". We have the following


"(a, b, c)=p(1,0,-1)+q(1,1,1)+r(1,0,0)=(p+q+r,q, -p+q)"


So, "a=p+q+r,\ \ b=q" and "c=-p+q" . Therefore, "p=q-c=b-c" and "r=a-p-q=a-(b-c)-b=a-2b+c" .


Therefore, "v = (a, b, c)=(b-c)a_1+ba_2+(a-2b+c)a_3."